# Predicting liquid properties and behavior via droplet pinch-off and machine learning

### v1

@@ Abstract
Here we demonstrate that the time-evolving interface observed during droplet formation, and consequently the resulting morphology nearing pinch-off, encode sufficient physical information for machine-learning (ML) frameworks to accurately infer key fluid properties, including viscosity and surface tension. Snapshots of dripping drops at the moment of break-up, together with their liquid properties and the flow rate, are used to form a data set for training ML algorithms. Experiments consisted of visualizing, using high-speed imaging, the process of droplet formation and identifying the frame closest to break-up. Experiments were conducted using Newtonian fluids under controlled flow conditions. In terms of the Reynolds (Re) and Ohnesorge (Oh) numbers, our conditions cover the domains 0.001 << Re << 200 and 0.01 << Oh << 20, by using silicon oils, aqueous solutions of ethanol and glycerin, and methanol. For each case, flow parameters were recorded, along with images capturing the final stages of droplet break-up. Supervised regression models were trained to predict fluid parameters from the extracted contours of the breaking droplets. Our data set contains 840 examples. The results demonstrate that the droplet geometry at pinch-off contains sufficient information to infer fluid properties by machine learning approaches. Our methods can predict surface tension, viscosity, or the droplet shape at pinch-off. These approaches provide alternatives to conventional methods to measure liquid properties while reducing measurement complexity and evaluation time and facilitating integration into automation. Unsupervised clustering is performed; the clusters represent regions in the Re–Oh and Bo–Oh planes, indicating that the latent representation may reveal physical properties and offering insight into droplet dynamics.
keywords:
Droplet $|$ Breakup $|$ Viscosity $|$ Surface tension $|$ Machine learning
@@
This manuscript was compiled on November 26, 2025
@@
The quantification of liquid properties, such as surface tension, viscosity, and density, is an essential quality-control measure in fluid dispensing processes such as inkjet printing(1, 2) and spray coating (3, 4). However, some conventional methods for determining fluid properties rely on specialized and complex instrumentation, which often limits their use within small enterprises and complicates automation. Among various existing techniques, the pendant drop method has become a widely adopted approach to determine the surface tension of a liquid (5). In this method, a droplet is suspended from the tip of a capillary, or needle, with the droplet contour recorded by optical imaging. The shape of the pendant droplet, and thus its contour, is controlled by a balance between gravity and interfacial forces (6). Once the droplet boundary is recorded, numerical solutions of the Young–Laplace equation, which accounts for the pressure difference across the liquid interface, the surface curvature, and surface tension (7), are fitted to determine the surface tension $\sigma$ (8). This method is valued for its simplicity, small liquid volume requirements, and compatibility with conventional optical imaging systems, making it suitable for both research and industrial environments (9). In terms of limitations, the pendant drop method relies on the accurate detection of the droplet boundary, which restricts its applicability in cases of low optical resolution, or high image noise, and/or for small droplets (10, 11, 12). Consequently, most setups work with millimetre-sized droplets (10, 12). Other experimental techniques used to characterise surface tension include the Wilhelmy plate, the Du Noüy ring, and the widely used bubble pressure tensiometry. In the Wilhelmy plate method the surface tension of a liquid is determined by measuring the force exerted by the liquid on a thin fully-wetted plate in contact with its surface (13). Alternatively, and in a similar fashion, the Du Noüy ring method works by lifting a thin metallic ring from a liquid surface while recording the force required to detach it (13). These two methods are simple, but their accuracy depends on the skills of the operator and on the use of materials that ensure the complete wetting of the plate or ring. On the other hand, bubble pressure tensiometers are reliable systems that obtain surface tension by measuring the pressure required to produce a bubble at the tip of a capillary submerged in the tested liquid (14). In this method, the bubble production rate is adjustable, making the technique particularly effective for studying time-dependent interfacial phenomena, such as surfactant diffusion and adsorption (15). However, bubble pressure tensiometers are usually expensive and less suitable for highly viscous fluids (16).
Viscosity is another key property that controls various aspects of fluid mechanics on the capillary scale, including jetting(17), spray dynamics(18), and liquid break-up (19). In industrial environments, viscosity is often measured by vibrational viscometers and rheometers. Vibrational methods work by quantifying the viscous damping of shear acoustic waves through the liquid. These viscometers are fast, easy to operate, and provide reliable results for liquids in which viscosity remains constant regardless of shear; these liquids are commonly referred to as ‘Newtonian fluids’. A disadvantage of vibrometers is that they require large volumes of liquid, often in the 200 ml range. Liquids with shear-dependent viscosity, such as shear-thinning or thickening, and viscoelastics, require capillary and/or rotational rheometers (20). These rheometers determine viscosity by applying a controlled shear stress, measuring the resulting flow response of the fluid. Unfortunately, rheometers are often both expensive and unsuitable for the fast-shear rates found in sprays and inkjet printing (21, 22).
Whilst effective, the traditional techniques described above form a segmented toolkit that lacks integration and are often incompatible with automation. In brief, in this work, we present an integrated, automation-friendly methodology, utilizing both supervised and unsupervised machine learning techniques, that can determine the surface tension, or the viscosity, or both, of Newtonian fluids from a single snapshot of a dripping drop. In our methods, the flow rate, the nozzle size, the density, and, if available, the viscosity or the surface tension are used as additional inputs.
In more detail, high-resolution and high-speed imaging experiments were conducted in this work to capture the last moments before the break-up of dripping droplets. From these images, droplet contours were extracted and used as key input features to supervised learning models. Various models are then trained to predict physical parameters, such as the viscosity and surface tension, directly from the image data. In addition, models are trained to predict the break-up shape of liquids given their material properties and flow rate. Unsupervised clustering algorithms are additionally employed to discover latent structures within the dataset, enabling the identification of distinct droplet behaviours, or quality modes, without requiring predefined labels. ML methods require sufficient and representative datasets, along with well-defined input features and target outputs; here, the training is carried out directly from experimental data, taken under a large variety of conditions, ensuring that the learned models capture the intrinsic complexity of wide laboratory conditions. This dual (supervised plus unsupervised) approach not only offers a fast and automated alternative to conventional rheometry and tensiometry but also reveals interpretable morphological patterns linked to fluid properties and experimental conditions. Ultimately, our integration of physics-informed (our hypothesis) imaging and machine learning offers a scalable and generalizable pathway for data-driven fluid diagnostics.
For over a century, the break-up of liquid jets and filaments has been at the forefront of fluid mechanics, given its relevance to a myriad of processes ranging from lava ejection and fragmentation to inkjet printing (23, 24). This phenomenon, while seemingly simple, involves complex dynamics where a liquid thins, or necks, rapidly until pinch-off, evolving through dramatic interfacial changes, all within a few microseconds (25). Indeed, the point of break-up is a finite-time singularity and its description requires both scaling arguments and high-resolution experiments and modelling to resolve features that arise in its vicinity. A comprehensive theoretical framework has classified the routes to break-up into various canonical regimes depending on the relative importance of inertia, capillarity, and viscosity (25).In the inviscid regime, where inertia balances surface tension, the thinning of a liquid neck follows the power law $h_{\text{min}}\sim(t_{b}-t)^{2/3}$, where $h_{\text{min}}$ is the minimum neck radius, $t_{b}$ is the break-up time and $t$ is time. Interestingly, liquid break-up under this regime shows self-similarity where, regardless of the flow conditions, the liquid interface near pinch-off converges toward a unique shape, offering a powerful criterion for identifying inviscid liquids(26, 27).This regime is often observed in low–viscosity fluids such as water or ethanol, when surrounded by a negligible ambient environment(28, 29).
In most other liquids, both viscosity and inertia play a dominant role. In fact, when viscous forces dominate, the thinning is characterized by a linear scaling, i.e. $h_{\min}=0.0709\left(\frac{\mathrm{\sigma}}{\mathrm{\mu}}\right)(t_{b}-t)$, where $\sigma$ is surface tension and $\mu$ is the viscosity(30). In contrast, in cases where all three forces, i.e. capillary, viscous, and inertial, play a role in the dynamics, the viscous–inertial regime leads to the universal scaling $h_{\min}=0.0304\left(\frac{\mathrm{\sigma}}{\mathrm{\mu}}\right)(t_{b}-t)$(31). A further fourth regime exists where the surrounding media viscosity, often another liquid, dictates the thinning and pinch-off dynamics(32). It is worth noting that, on the route to break-up, the neck dynamics transition between these regimes based on the local importance of inertia, viscosity and surface tension effects(19). The thinning regimes, and their transitions, are parametrized by two dimensionless parameters, the Reynolds ($Re=\frac{4\rho Q}{\pi\mu D}$), and the Ohnesorge ($Oh=\frac{\mu}{\sqrt{\rho\sigma D}}$) numbers, where $Q$ is the volumetric flow rate, $\rho$ is the liquid density, and $D$ is a characteristic length (often the nozzle size or the the neck diameter). In fact, the transitional behavior has been found to consist of multiple intermediate transient regimes, which are highly sensitive to the liquid and flow properties. In other words, in their route to break-up, two liquids of similar characteristics can have a very different thinning transitional behavior (19). Thus, the profile of a liquid neck near, or at, break-up is a repository of fluid dynamics information, reflecting the transitional competition of fluid properties and the flow at the neck. Complementary works have demonstrated that the break-up of liquid filaments, and the transition from dripping to jetting, are also ruled by the Reynolds and Ohnesorge numbers and by the filament length to diameter aspect ratio(33, 34). A further dimensionless parameter, in addition to $Re$ and $Oh$, the Bond number, $Bo=\frac{\Delta\rho gD_{d}^{2}}{\sigma}$, has been used to parametrize dripping from jetting, where $D_{d}$ denotes the droplet diameter after break-up. These past findings support our hypothesis that the route to break-up is unique for each liquid and thus the resulting shape at the time of pinch-off is unique. Under this hypothesis, we have created a large database of images showing the last moment before break-up of fully characterized dripping droplets, to train ML models and predict liquid properties.
Refer to captiona)
Refer to captionb)                       c)

### v2

# Predicting liquid properties and behaviour via droplet pinch-off and machine learning

@@ Abstract
Here we demonstrate that the time-evolving interface observed during droplet formation, and consequently the resulting morphology nearing pinch-off, encode sufficient physical information for machine-learning (ML) frameworks to accurately infer key fluid properties, including viscosity and surface tension. Snapshots of dripping drops at the moment of break-up, together with their liquid properties and the flow rate, are used to form a data set for training ML algorithms. Experiments consisted of visualizing, using high-speed imaging, the process of droplet formation by dripping, and identifying the frame closest to break-up. Experiments were conducted using Newtonian fluids under controlled flow conditions. In terms of the Reynolds (Re) and Ohnesorge (Oh) numbers, our conditions cover the domains 0.001 << Re << 200 and 0.01 << Oh << 20, by using silicon oils, aqueous solutions of ethanol and glycerin, and methanol. For each case, flow parameters were recorded, along with images capturing the final stages of droplet break-up. Supervised regression models were trained to predict fluid parameters from the extracted contours of the breaking droplets. Our data set contains 840 examples. Our results demonstrate that machine-learning models can accurately infer, across a wide range of fluids, key material properties using a single high-speed snapshot taken just before pinch-off. Our methods can predict surface tension, viscosity, or the droplet shape at pinch-off. These approaches provide alternatives to conventional methods to measure liquid properties while reducing measurement complexity and evaluation time and facilitating integration into automation. In addition, unsupervised clustering is performed; the clusters represent regions in the Re–Oh and Bo–Oh planes, indicating that the latent representation may reveal physical properties and offering insight into droplet dynamics.
I Introduction
The quantification of liquid properties, such as surface tension, viscosity, and density, is an essential quality-control measure in fluid dispensing processes such as inkjet printing[1, 2] and spray coating [3, 4]. However, some conventional methods for determining fluid properties rely on specialized and complex instrumentation, which often limits their use within small enterprises and complicates automation. Among various existing techniques, the pendant drop method has become a widely adopted approach to determine the surface tension of a liquid [5]. In this method, a droplet is suspended from the tip of a capillary, or needle, with the droplet contour recorded by optical imaging. The shape of the pendant droplet, and thus its contour, is controlled by a balance between gravity and interfacial forces [6]. Once the droplet boundary is recorded, numerical solutions of the Young–Laplace equation, which accounts for the pressure difference across the liquid interface, the surface curvature, and surface tension [7], are fitted to determine the surface tension $\sigma$ [8]. This method is valued for its simplicity, small liquid volume requirements, and compatibility with conventional optical imaging systems, making it suitable for both research and industrial environments [9]. In terms of limitations, the pendant drop method relies on the accurate detection of the droplet boundary, which restricts its applicability in cases of low optical resolution, or high image noise, and/or for small droplets [10, 11, 12]. Consequently, most setups work with millimetre-sized droplets [10, 12]. Other experimental techniques used to characterise surface tension include the Wilhelmy plate, the Du Noüy ring, and the widely used bubble pressure tensiometry. In the Wilhelmy plate method the surface tension of a liquid is determined by measuring the force exerted by the liquid on a thin fully-wetted plate in contact with its surface [13]. Alternatively, and in a similar fashion, the Du Noüy ring method works by lifting a thin metallic ring from a liquid surface while recording the force required to detach it [13]. These two methods are simple, but their accuracy depends on the skills of the operator and on the use of materials that ensure the complete wetting of the plate or ring. On the other hand, bubble pressure tensiometers are reliable systems that obtain surface tension by measuring the pressure required to produce a bubble at the tip of a capillary submerged in the tested liquid [14]. In this method, the bubble production rate is adjustable, making the technique particularly effective for studying time-dependent interfacial phenomena, such as surfactant diffusion and adsorption [15]. However, bubble pressure tensiometers are usually expensive and less suitable for highly viscous fluids [16].
Viscosity is another key property that controls various aspects of fluid mechanics on the capillary scale, including jetting[17], spray dynamics[18], and liquid break-up [19]. In industrial environments, viscosity is often measured by vibrational viscometers and rheometers. Vibrational methods work by quantifying the viscous damping of shear acoustic waves through the liquid. These viscometers are fast, easy to operate, and provide reliable results for liquids in which viscosity remains constant regardless of shear; these liquids are commonly referred to as ‘Newtonian fluids’. A disadvantage of vibrometers is that they require large volumes of liquid, often in the 200 ml range. Liquids with shear-dependent viscosity, such as shear-thinning or thickening, and viscoelastics, require capillary and/or rotational rheometers [20]. These rheometers determine viscosity by applying a controlled shear stress, measuring the resulting flow response of the fluid. Unfortunately, rheometers are often both expensive and unsuitable for the fast-shear rates found in sprays and inkjet printing [21, 22].
Whilst effective, the traditional techniques described above form a segmented toolkit that lacks integration and are often incompatible with automation. In this work, we present an integrated, automation-friendly methodology, utilizing both supervised and unsupervised machine learning techniques, that can determine the surface tension, or the viscosity, or both, of Newtonian fluids from a single snapshot of a dripping drop. In our methods, the flow rate, the nozzle size, the density, and, if available, the viscosity or the surface tension are used as additional inputs.
In more detail, high-resolution and high-speed imaging experiments were conducted in this work to capture the last moments before the break-up of dripping droplets. From these images, droplet contours were extracted and used as key input features to supervised learning models. Various models are then trained to predict physical parameters, such as the viscosity and surface tension, directly from the image data. In addition, models are trained to predict the break-up shape of liquids given their material properties and flow rate. Unsupervised clustering algorithms are additionally employed to discover latent structures within the dataset, enabling the identification of distinct droplet behaviours, or quality modes, without requiring predefined labels. ML methods require sufficient and representative datasets, along with well-defined input features and target outputs; here, the training is carried out directly from experimental data, taken under a large variety of conditions, ensuring that the learned models capture the intrinsic complexity of wide laboratory conditions. This dual (supervised plus unsupervised) approach not only offers a fast and automated alternative to conventional rheometry and tensiometry but also reveals interpretable morphological patterns linked to fluid properties and experimental conditions. Ultimately, our integration of physics-informed imaging (based on our hypothesis) and machine learning offers a scalable and generalizable pathway for data-driven fluid diagnostics.
II Background
II.1 Liquid breakup
For over a century, the break-up of liquid jets and filaments has been at the forefront of fluid mechanics, given its relevance to a myriad of processes ranging from lava ejection and fragmentation to inkjet printing [23, 24]. This phenomenon, while seemingly simple, involves complex dynamics where a liquid thins, or necks, rapidly until pinch-off, evolving through dramatic interfacial changes, all within a few microseconds [25]. Indeed, the point of break-up is a finite-time singularity and its description requires both scaling arguments and high-resolution experiments and modelling to resolve features that arise in its vicinity. A comprehensive theoretical framework has classified the routes to break-up into various canonical regimes depending on the relative importance of inertia, capillarity, and viscosity [25].In the inviscid regime, where inertia balances surface tension, the thinning of a liquid neck follows the power law $h_{\text{min}}\sim(t_{b}-t)^{2/3}$, where $h_{\text{min}}$ is the minimum neck radius, $t_{b}$ is the break-up time and $t$ is time. Interestingly, liquid break-up under this regime shows self-similarity where, regardless of the flow conditions, the liquid interface near pinch-off converges toward a unique shape, offering a powerful criterion for identifying inviscid liquids[26, 27].This regime is often observed in low–viscosity fluids such as water or ethanol, when surrounded by a negligible ambient environment[28, 29].
In most other liquids, both viscosity and inertia play a dominant role. In fact, when viscous forces dominate, the thinning is characterized by a linear scaling, i.e. $h_{\min}=0.0709\left(\frac{\mathrm{\sigma}}{\mathrm{\mu}}\right)(t_{b}-t)$, where $\sigma$ is surface tension and $\mu$ is the viscosity[30]. In contrast, in cases where all three forces, i.e. capillary, viscous, and inertial, play a role in the dynamics, the viscous–inertial regime leads to the universal scaling $h_{\min}=0.0304\left(\frac{\mathrm{\sigma}}{\mathrm{\mu}}\right)(t_{b}-t)$[31]. A further fourth regime exists where the surrounding media viscosity, often another liquid, dictates the thinning and pinch-off dynamics[32]. It is worth noting that, on the route to break-up, the neck dynamics transition between these regimes based on the local importance of inertia, viscosity and surface tension effects[19]. The thinning regimes, and their transitions, are parametrized by two dimensionless parameters, the Reynolds ($Re=\frac{4\rho Q}{\pi\mu D}$), and the Ohnesorge ($Oh=\frac{\mu}{\sqrt{\rho\sigma D}}$) numbers, where $Q$ is the volumetric flow rate, $\rho$ is the liquid density, and $D$ is a characteristic length (often the nozzle size or the neck diameter). In fact, the transitional behavior has been found to consist of multiple intermediate transient regimes, which are highly sensitive to the liquid and flow properties. In other words, in their route to break-up, two liquids of similar characteristics can have a very different thinning transitional behavior [19]. Thus, the profile of a liquid neck near, or at, break-up is a repository of fluid dynamics information, reflecting the transitional competition of fluid properties and the flow at the neck. Complementary works have demonstrated that the break-up of liquid filaments, and the transition from dripping to jetting, are also ruled by the Reynolds and Ohnesorge numbers and by the filament length to diameter aspect ratio[33, 34]. A further dimensionless parameter, in addition to $Re$ and $Oh$, the Bond number, $Bo=\frac{\Delta\rho gD_{d}^{2}}{\sigma}$, has been used to parametrize dripping from jetting, where $D_{d}$ denotes the droplet diameter after break-up. These past findings support our hypothesis that the route to break-up is unique for each liquid and thus the resulting shape at the time of pinch-off is unique. Under this hypothesis, we have created a large database of images showing the last moment before break-up of fully characterized dripping droplets, to train ML models and predict liquid properties.
II.2 Machine learning
Machine learning (ML) is a branch of computer modeling that learns patterns from large amounts of data and makes predictions, or decisions, without prior knowledge beyond what it has learned from this data. In recent years, ML has been applied across various physics topics, including fluid dynamics[35, 36, 37], droplet physics [38, 39, 40, 41, 42], soft matter [43, 44],and materials science [45]. These studies highlight the potential of ML as a tool to complement traditional experimental and theoretical approaches in science. Specifically, supervised learning trains models on labelled datasets (data that has been manually classified or characterised), mapping inputs to known outputs, which is particularly effective for prediction tasks. In contrast, unsupervised learning operates on unlabelled data to identify hidden patterns or underlying structures, i.e. clusters. Both strategies offer valuable means of extracting knowledge from complex experimental observations. In addition, recent advances in computer vision[46] provide powerful tools for analysing image-based experimental data [47], which are highly relevant for the study of droplet dynamics and pinch-off processes where high-speed imaging is often used in conjunction with image analysis to extract information from images[19].
II.3 ML Applications to Droplet Dynamics
ML methods have been applied to various systems involving droplet formation, examples of these are on direct inkjet writing[48], droplet impact, spreading and bouncing on nanostructure substrates[49], spreading on various materials [50], and droplet splashing [51]. Considering first the forward-prediction paradigm, in which ML models predict droplet behaviour from known material and operating conditions, Kim et al.[52] trained supervised regressors to predict the jetting behaviour of viscoelastic inks from rheological and drive parameters, with later work [53] using reinforcement learning to optimise actuation waveforms, dynamically improving jet stability without an explicit physical model. Additionally, Chen et al.[39] employed supervised learning to predict the optimum printable biomaterial formulations for direct ink writing, and Li et al.[41] developed a multi-parameter droplet optimisation model including flow and material parameters. Turning to image-based studies, Kratz and Kierfeld [40] introduced an image-based ML approach for pendant-drop tensiometry, demonstrating that the contour of a static droplet can be used to infer surface tension with the method restricted to steady pendant configurations. Pierzyna et al. [51] developed a data-driven model for the splashing thresholds and Au-Yeung et al. [49] developed a method to predict the maximum spreading and impact morphology of droplets on nanostructured surfaces.
In summary, previous works have predominantly used ML to control, or classify, droplet behaviour given known physical inputs. In contrast, the present work uses supervised learning to tackle the inverse problem of inferring fluid properties directly from droplet morphology, in addition, we use unsupervised learning to gain insight into the underlying physics of this process.
III Experimental Methodology
@@
Figure 1: Experimental setup used to capture the necking and break-up dynamics during droplet dripping. The system is a simple shadowgraph, where visualisation was carried out through a macro lens and a diffused light source.
@@
Liquid $\mu$ (mPa s) $\rho$ (kg/m3) $\sigma$ (mN/m)
Water 1.0 1,000 72.0
90% water / 10% glycerin 1.5 1,005 71.5
70% water / 30% glycerin 2.8 1,033 69.2
50% water / 50% glycerin 6.8 1,082 67.4
30% water / 70% glycerin 32.9 1,156 65.8
10% water / 90% glycerin 225.1 1,206 63.8
Glycerin 1,350.0 1,248 62.8
90% water / 10% ethanol 1.4 951 50.4
70% water / 30% ethanol 2.5 927 34.2
50% water / 50% ethanol 2.5 886 28.6
10% water / 90% ethanol 1.6 811 23.9
Ethanol 1.2 789 22.2
Silicon oil (2 cSt) 1.8 900 17.7
Silicon oil (50 cSt) 48.6 960 20.8
Silicon oil (350 cSt) 348.0 969 21.1
Silicon oil (1000 cSt) 996.4 970 21.2
Table 1: Physical properties of the tested liquids at $T=20.0\pm 1.0\,^{\circ}$C. Values of viscosity $\mu$, density $\rho$, and surface tension $\sigma$ used in this work.
III.1 Liquid materials and characterisation
[…]
Figure 2: Examples of the image-processing pipeline. The left panel shows the droplet at breakup, the middle panel shows the binary segmentation, and the right panel shows the extracted interface silhouette.
The physical properties, i.e. density, viscosity, and surface tension, of the test liquids were characterised at (working) room temperature of (21.0 $\pm$ 1.0 Celsius). Density was determined gravimetrically by weighing a known volume of liquid using an analytical balance. Viscosity was measured on a rotational rheometer (Anton Paar MCR 102e) equipped with a cone–plate geometry at a constant shear rate of $50\,\text{s}^{-1}$. Each sample was measured twenty times with the mean value taken as the working viscosity; the typical error in the viscosity measurement is $\delta\mu_{true}=\pm$ 0.10 mPa s. For liquids of low and moderate viscosity ( $\mu\lesssim 60~\mathrm{mPa\,s}$), the surface tension was evaluated via the maximum bubble pressure method using a SITA pro line T15 tensiometer at bubble lifetimes from 1.0 s to 15.0 s. Within this range, surface tension was found to be independent of surface age, with a standard error of $\delta\sigma_{true}=\pm 0.2$ mN/m. As noted in the introduction, bubble-pressure tensiometers are known to provide unreliable readings for highly viscous liquids due to viscous and hydrodynamic effects[16, 54]. Accordingly, for glycerin and water mixtures of a concentration $\geq 80$ wt% glycerin, and for 350 cSt and 1000 cSt silicone oils, surface tension values were taken from the manufacturers’ data-sheets. The resulting liquid property data is presented in Table 1; we note that viscosity values span more than three orders of magnitude while surface tension triples in value. The domain of our data, parametrized in terms of $Re$ and $Oh$, can be found in Fig. 3a. As can be seen there, our conditions cover four orders of magnitude in the Ohnesorge number and five in Reynolds, and include simple dripping, and jetting as defined in previous works [33].
III.2 Image Analysis
High-speed images were visually analysed to identify the nearest snapshot before break-up; this was done for each experiment. We carried out 840 conditions/experiments, each producing an image (snapshot) showing the break-up of the droplet. Each image was processed by custom MATLAB code to obtain the droplet profile and its contour. The code works by first subtracting a pre-recorded background (image without a droplet, just the background) from the droplet snapshot to enhance the interfacial contrast. The code then employs a global threshold function (Otsu’s method) to obtain a binary image[55]. This is followed by steps to remove image artefacts and holes. The largest connected region in the binary image is then identified as the droplet, and its exterior boundary is extracted as the raw contour. Finally, the boundary is smoothed and resampled uniformly. The needle outer diameter was used as size reference. Using this calibration, lengths, such as the droplet diameter $D_{d}$ at pinch-off, could be calculated and recorded for each experiment for further analysis. Figure 2 illustrates some of the steps of the image analysis, and the droplet shape at the point of break-up, for: a) pure glycerin; and b) water. These boundary-only silhouettes constitute the sole image inputs for our computational methods.